Assume the money supply is $850 billion, total deposits are $500 billion and the required reserve-deposit ratio is 10%, if the Central Bank purchases $60 million worth of Treasury bills, what is the greatest amount by which total money supply could change? Do you expect that money supply would actually change by that much? Find the maximum value of deposit multiplier for this economy. Explain, why this value is the maximum value.
Answer #1

Deposit multiplier = 1/required reserve ratio = 1/10%

Deposit multiplier = 10

If 60 million treasury bills are purchased,

Maximum Change in money supply = 60*10 = $600 million           (money supply will increase)

Money supply will not change up to that level, because some of the cash withdrawn, will not be deposited back in the bank. Some of the money will be kept at home for some emergency purposes. So, multiplier effect will not work and total maximization in money supply will not increase.